1.     Complete the following table.

(i)    Find out the total product (TP), average product (AP), marginal product (MP) and draw a graph of TP, AP, MP and also explain the stage of production.

(ii)  if the fixed price of land is 1000 and firm pay 750 to each worker then calculate, Fixed cost, variable Cost, Marginal Cost, Average variable Cost (AVC), Average Fixed Cost(AFC), and Average total cost(ATC).

Labor

Total Product

Average Product

Marginal Product

0

0

 

 

1

8

 

 

2

18

 

 

3

25

 

 

4

30

 

 

5

33

 

 

6

34

 

 



1
Expert's answer
2021-05-12T16:06:38-0400

(i)




Average Product "=\dfrac{\text{ Total Product}}{\text{ Labor}}"


Marginal Product"=\dfrac{\text{ change in total product}}{\text{ change in labor

}}"


The graph is-




The total product curve will always increase. It increases at an increasing rate then increases at a decreasing rate. And in the beginning AP curve is above MP curve but as labor increases MP goes above AP and MP interest AP at the minimum.



(ii)



FC is a fixed cost that does not change as labor changes.


VC is variable cost that changes with each labor


TC = VC + FC


"ATC = \dfrac{\text{TC}}{ \text{Labour}}"


"AVC =\dfrac{\text{VC} }{\text{ labor}}
\\[9pt]


AFC = \dfrac{ \text{FC }}{ Labor}

\\[9pt]

MC = \dfrac{\text{change in TC }}{ \text{change in labor}}"


The fixed cost, variable costs, average total cost , average variable cost are shown in the above table.